# and Qian Liu

### v1

@@ II-B Tri-hybrid Beamforming Architecture
BB domain: We exploit the digital beamformer $\mathbf{F}_{\mathrm{BB}}\triangleq[\mathbf{f}_{\mathrm{BB},1},\mathbf{f}_{\mathrm{BB},2},\ldots,\mathbf{f}_{\mathrm{BB},K}]\in\mathbb{C}^{N_{\mathrm{RF}}\times K}$ to pre-process transmitted signals in the BB domain, addressing the limitation of having fewer RF chains, which restricts direct control over each antenna’s signal. This procedure effectively mitigates multi-user interference and enhances overall system performance.
@@ II-C Channel Model
After establishing the mathematical formulation of the tri-hybrid beamforming architecture, our focus now shifts to developing a comprehensive channel model that accommodates the expanded dimensionality introduced by the pattern-reconfigurable antenna array. Specifically, unlike conventional channel models that primarily account for path loss and spatial phase gradients, the proposed model incorporates the virtual angular index channel matrix $\overline{\mathbf{H}}_{\mathrm{EM},k},\forall k$ in (2). This integration enables an explicit representation of the angular characteristics of each propagation path. Furthermore, when combined with the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in (1), it enables the determination of the radiation intensity $\mathbf{G}_{\mathrm{EM},k},\forall k$ in (3), facilitating a more precise characterization of the system’s electromagnetic behavior.
To establish a foundation for the innovative channel model of the pattern-reconfigurable antenna array, we first present the traditional mmWave channel model. Due to the sparse scattering characteristics of the mmWave band, the channel typically consists of a few dominant clusters, each containing multiple propagation paths [28]. Let $\mathcal{C}\triangleq\{1,2,\ldots,C\}$ denote the set of scattering clusters, and $\mathcal{L}_{c}\triangleq\{1,2,\ldots,L_{c}\}$ denote the set of multi-path rays within the $c$-th scattering cluster. The channel between the $n$-th antenna and the $k$-th UE can be expressed as
$h_{k,n}=\sum_{c=1}^{C}\sum_{l_{c}=1}^{L_{c}}\alpha_{k,c,l_{c}}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,c,l_{c}}},~{}\forall k,n,$ (4)
where $\theta_{k,c,l_{c}}$ and $\alpha_{k,c,l_{c}}$ denote the angle of departure (AoD) and channel gain corresponding to the $l_{c}$-th ray in the $c$-th scattering cluster, respectively. The AoD $\theta_{k,c,l_{c}}$ can be described as
@@
where $\overline{\theta}_{k,c}$ denotes the nominal AoD of the $c$-th scattering cluster for the $k$-th UE, and $\Delta\theta_{k,l_{c}}$ denotes the angular deviation of the $l_{c}$-th ray from the nominal direction, following Gaussian distribution $\Delta\theta_{k,l_{c}}\sim\mathcal{N}(0,\varsigma^{2}_{k,c})$, with $\varsigma_{k,c}$ indicating the angle spread. The channel gain involves the typically distance-dependent path loss with $r_{k,c,l_{c}}$ representing the propagation distance. It follows the same modeling as in (5), i.e., $r_{k,c,l_{c}}=\overline{r}_{k,c}+\Delta r_{k,l_{c}}$, where $\Delta r_{k,l_{c}}\sim\mathcal{N}(0,\iota^{2}_{k,c})$. For better clarity in the subsequent channel model formulation, we define $L\triangleq\sum_{c=1}^{C}L_{c}$ as the total number of paths. With this notation, the channel model in (4) can be reformulated in a more compact form as follows
$h_{k,n}=\sum_{l=1}^{L}\alpha_{k,l}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}},~{}\forall k,n.$ (6)
@@
Furthermore, the integration of pattern-reconfigurable antennas extends the channel representation from a one-dimensional to an $M$-dimensional formulation by incorporating additional EM domain channel information, i.e., $\overline{\mathbf{H}}_{\mathrm{EM},k},\forall k$ in (2). Specifically, the dimension-extended channel for the $n$-th pattern-reconfigurable antenna, denoted as $\mathbf{h}_{k,n}\in\mathbb{C}^{M}$, is expressed as follows
@@
$\mathbf{h}_{k,n}=\sum_{l=1}^{L}(\alpha_{k,l}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}})\otimes\mathbf{h}_{\mathrm{EM},k,l},~{}\forall k,n.$ (7)
@@
Following this formulation, by stacking the channel vectors $\mathbf{h}_{k,n},\forall n$ of all $N_{\mathrm{t}}$ antennas, we obtain the complete channel vector for the $k$-th UE, given by $\mathbf{h}_{k}\triangleq[\mathbf{h}_{k,1}^{T},\mathbf{h}_{k,2}^{T},\ldots,\mathbf{h}_{k,N_{\mathrm{t}}}^{T}]^{T}\in\mathbb{C}^{MN_{\mathrm{t}}}$. The corresponding expression can be explicitly formulated as
@@
$\mathbf{h}_{k}=\sum_{l=1}^{L}(\alpha_{k,l}\mathbf{a}_{k,l})\otimes\mathbf{h}_{\mathrm{EM},k,l},~{}\forall k,$ (8)
where
@@
$\mathbf{a}_{k,l}$ $\triangleq[1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\theta_{k,l}},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\theta_{k,l}}]^{T},~{}\forall k,l,$ (9)
denotes the steering vector. By applying the Kronecker product formula and rearranging the order of summation, (8) can be equivalently expressed as
$\mathbf{h}_{k}=\mathrm{vec}(\mathbf{H}_{\mathrm{EM},k}\mathbf{H}_{\mathrm{S},k}^{T}),~{}\forall k,$ (10)
with
$\mathbf{H}_{\mathrm{S},k}$ $=[\alpha_{k,1}\mathbf{a}_{k,1},\alpha_{k,2}\mathbf{a}_{k,2},\ldots,\alpha_{k,L}\mathbf{a}_{k,L}]\in\mathbb{C}^{N_{\mathrm{t}}\times L},~{}\forall k,$ (11)
representing the spatial domain channel. Additionally, to enhance clarity and brevity, the EM domain channel and spatial domain channel can be separated into two independent multiplicative expressions, which yields
$\mathbf{h}_{k}=\overline{\mathbf{H}}_{\mathrm{EM},k}\overline{\mathbf{h}}_{\mathrm{S},k},~{}\forall k,$ (12)
@@
where $\overline{\mathbf{H}}_{\mathrm{EM},k}\in\{0,1\}^{MN_{\mathrm{t}}\times LN_{\mathrm{t}}}$ is defined in (2) containing the EM domain channel information, and $\overline{\mathbf{h}}_{\mathrm{S},k}\triangleq\mathrm{vec}(\mathbf{H}_{\mathrm{S},k}^{T})\in\mathbb{C}^{LN_{\mathrm{t}}}$ includes the spatial domain channel information.
@@ II-D System Model
$y_{k}$ $=\overline{\mathbf{h}}_{\mathrm{S},k}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{x}+n_{k},~{}\forall k,$ (13)
@@
$\mathrm{SINR}_{k}=\frac{|\mathbf{\overline{h}}_{\mathrm{S},k}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}|\mathbf{\overline{h}}_{\mathrm{S},k}^{H}\mathbf{\overline{H}}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}},~{}\forall k,\$ (14)
@@ II-E Problem Formulation
In this paper, based on the tri-hybrid beamforming architecture, we propose to jointly design the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain, and the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain to maximize the sum-rate, while satisfying the constraints of radiation pattern energy, constant modulus and discrete phases of PSs, non-overlapping dynamic connection network, and total transmit power budget. Thus, the joint three domain beamformer optimization problem with full instantaneous CSI is formulated as
@@
$\mathrm{s.t.}$ $\ \ \|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,~{}\forall n,$ (16b)
@@
Within this tri-hybrid beamforming architecture, accurately acquiring instantaneous CSI is a significant challenge. The analog beamformer in the RF domain suffers from inherent undersampling, while the need to estimate instantaneous CSI for all possible radiation patterns in the EM domain leads to excessive pilot overhead. Moreover, the real-time computation of beamforming matrices across all three domains imposes high computational complexity, severely impacting system processing efficiency. To address these issues, we propose an innovative tri-timescale framework tailored for tri-hybrid beamforming design. This framework decomposes the original non-convex optimization problem in (16) into three independent subproblems, each solvable at a different timescale. This decomposition significantly reduces pilot overhead, enhances computational tractability, and improves overall system efficiency.
@@ III-A Tri-timescale Framework
The tri-hybrid beamforming architecture presents significant practical challenges, including high channel estimation overhead and computational complexity. Estimating the full CSI and jointly optimizing beamformers across all three domains in every time-slot is computationally infeasible. To address this, we propose a tri-timescale beamforming framework that effectively balances practical feasibility and system performance, as illustrated in Fig. 4.
The overall time domain corresponds to the channel statistics coherence time, during which the statistical distribution of the channel remains stable. This period is structured hierarchically, with beamforming operations executed across three distinct timescales, as detailed below:
  • Long-timescale: The total duration is divided into $T_{\mathrm{L}}$ super-frames, denoted as $\mathcal{T}_{\mathrm{L}}\triangleq\{1,2,\ldots,T_{\mathrm{L}}\}$. In each super-frame, the BS extracts high-dimensional channel information of the scattering cluster core to design the radiation beamformer, ensuring radiation pattern alignment in the EM domain.
  • Medium-timescale: Each super-frame consists of $T_{\mathrm{M}}$ frames, defined as $\mathcal{T}_{\mathrm{M}}\triangleq\{1,2,\ldots,T_{\mathrm{M}}\}$. In each frame, the BS acquires a statistical channel sample of multi-path rays within the scattering clusters to update the analog beamformer, enhancing array gain in the RF domain.
  • Short-timescale: Each frame is further divided into $T_{\mathrm{S}}$ time-slots, referred to as the channel coherence time, represented as $\mathcal{T}_{\mathrm{S}}\triangleq\{1,2,\ldots,T_{\mathrm{S}}\}$. In each time-slot, the BS obtains an instantaneous low-dimensional effective CSI to refine the digital beamformer, mitigating multi-user interference and accounting for fast channel fluctuations in the BB domain.

### v2

@@ II-B Tri-hybrid Beamforming Architecture
BB domain: The digital beamformer $\mathbf{F}_{\mathrm{BB}}$ operates at the baseband level to effectively suppress multi-user interference and optimize overall system performance. Specifically, it is defined as $\mathbf{F}_{\mathrm{BB}}\triangleq[\mathbf{f}_{\mathrm{BB},1},\mathbf{f}_{\mathrm{BB},2},\ldots,\mathbf{f}_{\mathrm{BB},K}]\in\mathbb{C}^{N_{\mathrm{RF}}\times K}$. By performing precise pre-processing of transmitted signals, the digital beamformer compensates for the limited number of RF chains, ensuring refined control over each user’s signal in the spatial domain and significantly enhancing system throughput.
@@ II-C Channel Model
After establishing the mathematical formulation of the tri-hybrid beamforming architecture, our focus now shifts to developing a comprehensive channel model that accommodates the expanded dimensionality introduced by the pattern-reconfigurable antenna array. Specifically, unlike conventional channel models that primarily account for path loss and spatial phase gradients, our proposed model explicitly integrates the virtual angular index channel matrix $\mathbf{H}_{\mathrm{EM},k,l}$, defined in (5), to represent the angular characteristics of each propagation path. By incorporating this channel matrix along with the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ described in (4), our approach precisely determines the radiation gain $\mathbf{G}_{\mathrm{EM},k,l}$ in (6), enabling a more accurate and comprehensive characterization of the EM domain behavior.
To establish a foundation for the innovative channel model of the pattern-reconfigurable antenna array, we first revisit the traditional mmWave channel model. Given the sparse scattering characteristics of the mmWave band, the channel typically consists of $C$ dominant scattering clusters, each containing $L_{c}$ multi-path rays [31], [32]. The channel between the $n$-th antenna and the $k$-th UE can be expressed as
$h_{k,n}=\sum_{c=1}^{C}\sum_{l_{c}=1}^{L_{c}}\alpha_{k,c}e^{j\phi_{k,l_{c}}}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,c,l_{c}}},~\forall k,n,$ (7)
where $\theta_{k,c,l_{c}}$ denotes the angle of departure (AoD) corresponding to the $l_{c}$-th ray in the $c$-th scattering cluster, respectively, $\alpha_{k,c}$ denotes the mean large-scale channel gain of the $c$-th scattering cluster, and $\phi_{k,l_{c}}$ denotes the small-scale phase term of the $l_{c}$-th ray. The AoD $\theta_{k,c,l_{c}}$ can be described as
@@
where $\overline{\theta}_{k,c}$ denotes the nominal AoD of the $c$-th scattering cluster for the $k$-th UE, and $\Delta\theta_{k,l_{c}}$ denotes the angular deviation of the $l_{c}$-th ray from the nominal direction, following Gaussian distribution $\Delta\theta_{k,l_{c}}\sim\mathcal{N}(0,\varsigma^{2}_{k,c})$, with $\varsigma_{k,c}$ indicating the angle spread. The mean large-scale channel gain $\alpha_{k,c}$ mainly involves the typically distance-dependent path loss with $r_{k,c}$ representing the propagation distance. The small-scale parameter $\phi_{k,l_{c}}$ follows a uniform distribution in $0,2\pi]$ and is independent from ray to ray. For better clarity in the subsequent channel model formulation, we define $L\triangleq\sum_{c=1}^{C}L_{c}$ as the total number of paths. With this notation, the channel model in (7) can be reformulated in a more compact form as
$h_{k,n}=\sum_{l=1}^{L}\alpha_{k,l}e^{\jmath\phi_{k,l}}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}},~\forall k,n.$ (9)
@@
Furthermore, the integration of pattern-reconfigurable antennas extends the channel representation from a one-dimensional to an $M$-dimensional formulation by incorporating additional EM domain channel information, i.e., $\mathbf{H}_{\mathrm{EM},k,l}$ in (5). Specifically, the dimension-extended channel for the $n$-th pattern-reconfigurable antenna $\mathbf{h}_{k,n}\in\mathbb{C}^{M}$ is expressed as
@@
$\mathbf{h}_{k,n}\!\!=\!\!\sum_{l=1}^{L}(\alpha_{k,l}e^{\jmath\phi_{k,l}}e^{-\jmath\frac{2\pi}{\lambda}(n-1)d\sin\theta_{k,l}}\!)\!\otimes\!\mathbf{h}_{\mathrm{EM},k,l},~\forall k,n.\!\!\!\!$ (10)
@@
Following this formulation, by stacking the channel vectors $\mathbf{h}_{k,n}$ of all antennas, we obtain the comprehensive channel vector for the $k$-th UE, given by $\mathbf{h}_{k}\triangleq[\mathbf{h}_{k,1}^{T},\mathbf{h}_{k,2}^{T},\ldots,\mathbf{h}_{k,N_{\mathrm{t}}}^{T}]^{T}\in\mathbb{C}^{MN_{\mathrm{t}}}$. The corresponding expression can be explicitly formulated as
$\mathbf{h}_{k}$ $=\sum_{l=1}^{L}(\alpha_{k,l}e^{\jmath\phi_{k,l}}\mathbf{a}_{k,l})\otimes\mathbf{h}_{\mathrm{EM},k,l},$ (11)
@@
$\triangleq\sum_{l=1}^{L}\underbrace{(\mathbf{I}_{N_{\mathrm{t}}}\otimes\mathbf{h}_{\mathrm{EM},k,l})}_{\mathbf{H}_{\mathrm{EM},k,l}}\underbrace{(\alpha_{k,l}e^{\jmath\phi_{k,l}}\mathbf{a}_{k,l})}_{\mathbf{h}_{\mathrm{S},k,l}},~\forall k,$
where the steering vector is defined as
@@
$\!\!\!\mathbf{a}_{k,l}$ $\!\triangleq\![1,e^{-\jmath\frac{2\pi}{\lambda}d\sin\theta_{k,l}},\ldots,e^{-\jmath\frac{2\pi}{\lambda}(N_{\mathrm{t}}-1)d\sin\theta_{k,l}}]^{T},~\forall k,l.\!\!\!\!\!\!\!\!$ (12)
The comprehensive channel $\mathbf{h}_{k}$ in (11) is composed of the virtual angular index channel matrix $\mathbf{H}_{\mathrm{EM},k,l}$ defined in (5) and the spatial domain channel $\mathbf{h}_{\mathrm{S},k,l}$ given by
$\mathbf{h}_{\mathrm{S},k,l}\triangleq\alpha_{k,l}e^{\jmath\phi_{k,l}}\mathbf{a}_{k,l},~\forall k,l.$ (13)
When the virtual angular index channel matrix $\mathbf{H}_{\mathrm{EM},k,l}$ interacts with the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in (4), the actual radiation gains of the antennas $\mathbf{G}_{\mathrm{EM},k,l}$ mentioned in (6) are obtained. These gains subsequently modify the conventional spatial domain channel $\mathbf{h}_{\mathrm{S},k,l}$, thereby shaping the effective channel seen by the analog and digital beamformers. Additionally, to enhance clarity and brevity, we replace the path-wise summation in (11) with an equivalent block-stacked multiplicative expressions, which yields
$\mathbf{h}_{k}\triangleq\mathbf{H}_{\mathrm{EM},k}\mathbf{h}_{\mathrm{S},k},~\forall k,$ (14)
@@
where $\mathbf{H}_{\mathrm{EM},k}\triangleq[\mathbf{H}_{\mathrm{EM},k,1},\mathbf{H}_{\mathrm{EM},k,2},\ldots,\mathbf{H}_{\mathrm{EM},k,L}]\in\{0,1\}^{MN_{\mathrm{t}}\times LN_{\mathrm{t}}}$ contains the virtual angular index channel matrices, and $\mathbf{h}_{\mathrm{S},k}\triangleq[\mathbf{h}_{\mathrm{S},k,1}^{T},\mathbf{h}_{\mathrm{S},k,2}^{T},\ldots,\mathbf{h}_{\mathrm{S},k,L}^{T}]^{T}\in\mathbb{C}^{LN_{\mathrm{t}}}$ includes the spatial domain channel information.
@@ II-D System Model
$y_{k}$ $=\mathbf{h}_{k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{x}+n_{k}$ (15)
$=\mathbf{h}_{\mathrm{S},k}^{H}\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{x}+n_{k},~\forall k,$
@@
$\mathrm{SINR}_{k}=\frac{|\mathbf{h}_{\mathrm{S},k}^{H}\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},k}|^{2}}{\sum_{j=1,j\neq k}^{K}|\mathbf{h}_{\mathrm{S},k}^{H}\mathbf{H}_{\mathrm{EM},k}^{H}\mathbf{F}_{\mathrm{EM}}\mathbf{F}_{\mathrm{RF}}\mathbf{f}_{\mathrm{BB},j}|^{2}+\sigma_{k}^{2}},~\forall k,\\$ (16)
@@ II-E Problem Formulation
In this paper, we propose a comprehensive tri-hybrid beamforming strategy, jointly optimizing the radiation beamformer $\mathbf{F}_{\mathrm{EM}}$ in the EM domain, the analog beamformer $\mathbf{F}_{\mathrm{RF}}$ in the RF domain, and the digital beamformer $\mathbf{F}_{\mathrm{BB}}$ in the BB domain. The primary objective is to maximize the sum-rate performance of the system while satisfying several critical constraints, including radiation pattern energy normalization, constant modulus and discrete-phase constraints of the PSs, non-overlapping dynamic subarray assignments, and a total transmit power budget. Therefore, the joint optimization problem encompassing the three beamforming domains under the assumption of perfect instantaneous CSI is formulated as
@@
$\mathrm{s.t.}\ \ \ \ \$$\ \ \|\mathbf{f}_{\mathrm{EM},n}\|^{2}=1,~\forall n,$ (18b)
@@
Accurately obtaining instantaneous CSI poses significant practical difficulties within this tri-hybrid framework. In the RF domain, the analog beamformer inherently suffers from undersampling due to the limited number of RF chains, while acquiring instantaneous CSI for all possible radiation patterns in the EM domain leads to excessive pilot overhead. In addition, real-time joint optimization of the beamforming matrices across all three domains results in high computational complexity and low processing efficiency. These challenges motivate the development of a more structured beamforming strategy, which will be addressed in the next section.
@@ III-A Tri-timescale Framework
In this section, we present a tri-timescale beamforming framework for the proposed tri-hybrid architecture. Instead of estimating full instantaneous CSI and jointly optimizing the radiation, analog, and digital beamformers in every time-slot, the design separates their updates across three different timescales so as to balance performance and implementation complexity, as illustrated in Fig. 4. The overall time horizon corresponds to the channel statistics coherence time, during which the statistical distribution of the channel remains approximately constant [^1] Our channel modeling follows the quasi-static clustered model described in 3GPP TR 38.901 [31], where cluster-level parameters (e.g., nominal angles and angle spreads) remain approximately constant within a channel-statistics coherence interval. This assumption is well suited for low-to-moderate mobility scenarios such as indoor hotspots and small-cell mmWave deployments. Extending the proposed tri-timescale framework to fully dynamic, high-mobility scenarios is an important topic for future research.. This period is organized hierarchically, with beamforming operations carried out over three distinct timescales, as detailed below:
  • Long-timescale: The total duration, corresponding to the channel-statistics coherence time, is divided into $T_{\mathrm{L}}$ super-frames, denoted as $\mathcal{T}_{\mathrm{L}}\triangleq\{1,2,\ldots,T_{\mathrm{L}}\}$. This coherence interval typically spans from a few seconds to several tens of seconds. In each super-frame, the BS extracts high-dimensional channel information of the scattering cluster core to design the radiation beamformer, ensuring radiation pattern alignment in the EM domain.
  • Medium-timescale: Each super-frame consists of $T_{\mathrm{M}}$ frames, denoted as $\mathcal{T}_{\mathrm{M}}\triangleq\{1,2,\ldots,T_{\mathrm{M}}\}$. The duration of each frame is determined by the sampling frequency of the statistical channel information, generally ranging from several hundred milliseconds to a few seconds. In each frame, the BS acquires a statistical channel sample of multi-path rays within the scattering clusters to update the analog beamformer, enhancing array gain in the RF domain.
  • Short-timescale: Each frame is further divided into $T_{\mathrm{S}}$ time-slots, denoted as $\mathcal{T}_{\mathrm{S}}\triangleq\{1,2,\ldots,T_{\mathrm{S}}\}$. Each time-slot is referred to as the channel coherence time, typically on the order of tens of milliseconds [^2] The durations of the three timescales depend on the carrier frequency, user mobility, and propagation environment. The orders of magnitude reported here follow typical mmWave deployments under the 3GPP TR 38.901 clustered channel model and serve only as representative examples [29]. In practice, the time-scale durations can be flexibly adjusted to match specific deployment scenarios without altering the proposed tri-timescale optimization framework.. In each time-slot, the BS obtains an instantaneous low-dimensional effective CSI to refine the digital beamformer, mitigating multi-user interference and accounting for fast channel fluctuations in the BB domain.